A horizontal pipeline is 360 mm in diameter and carries oil at a rate of 150 kg/s. This oil has a mass density of 850 kg/m3 and a viscosity of 0.224 Pa?sat the temperature of the pipeline.

If the pressure in the oil is 3.75 atm at the front end, how long can the pipeline be so that the pressure at the far end is no less than 2.50 atm? L = ???
Answer #1

8n0 The volumetric flow rate is, 17Q = 150/ 850 = 0.1764 m3/s


\Delta P = 3.75atm - 2.5 atm

= (3.75 - 2.5)*1.013*105

= 1.266*105 pa

L = \frac{\pi* 0.180^{4}}{8*0.224*0.176}*(1.266*10^{5} pa)

= 1.323*103m